OpenStudy (anonymous):

Pablo rolled a standard die sixty times. He got a 1 twelve times. How does this result compare to the expected results? Explain.

OpenStudy (cwrw238):

the probability of getting a 1 on one throw = 1/6

OpenStudy (cwrw238):

so the expected result for 12 throws = 1/6 * 12 = 2 so you should be able to answer your question now